### v2

@@ Appendix C Characterization of matrices with constant VC-dimension
Fact 1 provides a structural characterization of some matrices with constant VC-dimension, and it is an analogous result to the fact that non-trivial hereditary classes of graphs have low VC-dimension. To prove Fact 1, we introduce the notion of a bipartisation of a graph.
Definition C.1.
@@
Lemma C.2 (Lemma 3.4 of Bousquet et al. (2015)).
@@
Proof C.5.
Since $\mathbf{M}\in\mathcal{M}$ is Boolean, we can treat it as a bipartite graph $G$, with rows representing left vertices and columns representing right vertices. In particular, we can treat $\mathcal{M}$ as a class of bipartite graphs closed under vertex deletion (row deletion = deleting a left vertex, column deletion = deleting a right vertex). Thus we can apply lemma C.2 to $\mathcal{M}$.
@@
We conclude, that set of bipartite graphs corresponding to $\mathcal{M}$ does not contain any graph whose bipartisation is $H^{\prime}$. Hence by Corollary C.3 the VC-dimension must be a finite constant. Note that this is equivalent to the VC-dimension of the adjacency matrices being bounded, but $\mathcal{M}$ are not adjacency matrices. The adjacency matrices are of form
@@
for matrices $\mathbf{M}\in\mathcal{M}$. We already argued that each adjacency matrix has constant VC-dimension, and since removing rows does not increase the VC-dimension (it is equivalent to deleting sets), this implies $\mathbf{M}$ also has constant VC-dimension. In conclusion, each matrix in $\mathcal{M}$ has constant VC-dimension.
C.1 Examples of low VC-dimension matrices
@@
Eppstein (1995) showed that $H$-minor free directed graphs (where the minor can be obtained through vertex deletion, edge deletion, and edge contraction) has $\mathrm{VC}$-dimension at most $|V(H)|-1$. As an immediate consequence, since bipartite-graphs are triangle ($K_{3}$) free, the VC-dimension of any bipartite graph is $2$. As a consequence, given any adjacency matrix $\mathbf{A}$ of a $H$-minor free graph $G$, for any $k\geq 1$, $\mathbf{A}^{k}$ is the adjacency matrix of a $H$-minor free graph $G^{\prime}$ (since $\mathbf{A}^{k}$ corresponds to the reachability graph where an edge $(u,v)\in E(G^{\prime})$ corresponds to the existence of a $k$-length walk between $u$ and $v$ in $G$). Therefore, the VC-dimension of $\mathbf{A}^{k}$ is also at most $|V(H)|-1$.
Adjacency matrices of interval graphs.
@@
Interval graphs are the intersection graphs of a family of intervals on the real line. Then, since the hypothesis class of intervals on the real line =$\mathcal{H}=\{\mathbbm{1}_{a,b}:a\leq x\leq b\}$ has VC-dimension $2$ (any two points can be shattered by choosing an interval that includes one or both points, but no set of three points can always be shattered), every interval graph has $\mathrm{VC}$-dimension at most $2$.
Boolean kernel matrices.
@@
A matrix where each column $i$ (and each row) receives some label $\ell_{i}\in L$ from some (possibly infinite) set of possible labels $L$. Then, suppose each entry of $\mathbf{M}$ satisfies $\mathbf{M}_{i,j}=f(\ell_{i},\ell_{j})$ for some function $f:L\to\{0,1\}$. For example, the labels could be points in $L=\mathbb{R}^{d}$ with $f$ being indicator that the distance is at most some threshold. If $f,L$ is fixed (i.e., independent of the number of rows/columns) then these graphs are closed under row/column deletion and thus have bounded VC-dimension by Section 1.
Shortest-path structures
@@
Let $G$ be an undirected graph with non-negative edge weights and let $S$ be a subset of its shortest paths such that, for every pair $(u,v)$ of distinct vertices, $S$ contains exactly one shortest path between $u$ and $v$. Marie de Lima et al. (2023) defines a range space associated with $S$ and proves that its VC dimension is 2.
Adjacency matrices of planar graphs.
@@
Let $G$ be a planar graph. Then, there is a planar drawing of $G$, and deleting vertices and edges retains this property. Therefore, these graphs are closed under row/column deletion and thus have bounded VC-dimension by Section 1.
Adjacency matrices of semi-algebraic graphs of bounded description complexity.
@@
By the Milnor-Thom theorem in real algebraic geometry (Matoušek, 2002) and the Sauer-Shelah lemma (Assouad, 1983), semi-algebraic graphs of bounded description complexity have constant VC-dimensions. In turn, this extends to adjacency matrices of string graphs where every two curves intersect.
Practical matrices.
@@
Coudert et al. (2024) computes the VC-dimension of families of graphs that arise in practice, from protein interaction networks to autonomous Internet systems to web graphs. Even for such graphs with millions of nodes, the VC-dimension of the graphs are typically between $3$ and $8$.

### v3

@@ Appendix D Characterization of matrices with constant VC-dimension
Fact 1.1 provides a structural characterization of some matrices with constant VC-dimension, and it is an analogous result to the fact that non-trivial hereditary classes of graphs have low VC-dimension. To prove Fact 1.1, we introduce the notion of a bipartisation of a graph.
Definition D.1.
@@
Lemma D.2 (Lemma 3.4 of Bousquet et al. [2015]).
@@
Proof.
Since $\mathbf{M}\in\mathcal{M}$ is Boolean, we can treat it as a bipartite graph $G$, with rows representing left vertices and columns representing right vertices. In particular, we can treat $\mathcal{M}$ as a class of bipartite graphs closed under vertex deletion (row deletion = deleting a left vertex, column deletion = deleting a right vertex). Thus we can apply Lemma˜D.2 to $\mathcal{M}$.
@@
We conclude, that set of bipartite graphs corresponding to $\mathcal{M}$ does not contain any graph whose bipartisation is $H^{\prime}$. Hence by Corollary˜D.3 the VC-dimension must be a finite constant. Note that this is equivalent to the VC-dimension of the adjacency matrices being bounded, but $\mathcal{M}$ are not adjacency matrices. The adjacency matrices are of form
@@
for matrices $\mathbf{M}\in\mathcal{M}$. We already argued that each adjacency matrix has constant VC-dimension, and since removing rows does not increase the VC-dimension (it is equivalent to deleting sets), this implies $\mathbf{M}$ also has constant VC-dimension. In conclusion, each matrix in $\mathcal{M}$ has constant VC-dimension. ∎
D.1 Examples of low VC-dimension matrices
@@
Adjacency matrices of $H$-minor free graphs. Eppstein [1995] showed that $H$-minor free directed graphs (where the minor can be obtained through vertex deletion, edge deletion, and edge contraction) has $\mathrm{VC}$-dimension at most |V(H)|1|V(H)|-1. As an immediate consequence, since bipartite-graphs are triangle ($K_{3}$) free, the VC-dimension of any bipartite graph is $2$. As a consequence, given any adjacency matrix $\mathbf{A}$ of a $H$-minor free graph $G$, for any $k\geq 1$, $\mathbf{A}^{k}$ is the adjacency matrix of a $H$-minor free graph $G^{\prime}$ (since $\mathbf{A}^{k}$ corresponds to the reachability graph where an edge $(u,v)\in E(G^{\prime})$ corresponds to the existence of a $k$-length walk between $u$ and $v$ in $G$). Therefore, the VC-dimension of $\mathbf{A}^{k}$ is also at most |V(H)|1|V(H)|-1.
@@
Adjacency matrices of interval graphs. Interval graphs are the intersection graphs of a family of intervals on the real line. Then, since the hypothesis class of intervals on the real line $\mathcal{H}=\{\mathbbm{1}_{a,b}:a\leq x\leq b\}$ has VC-dimension $2$ (any two points can be shattered by choosing an interval that includes one or both points, but no set of three points can always be shattered), every interval graph has $\mathrm{VC}$-dimension at most $2$.
@@
Boolean kernel matrices. A matrix where each column $i$ (and each row) receives some label $\ell_{i}\in L$ from some (possibly infinite) set of possible labels $L$. Then, suppose each entry of $\mathbf{M}$ satisfies $\mathbf{M}_{i,j}=f(\ell_{i},\ell_{j})$ for some function $f:L\to\{0,1\}$. For example, the labels could be points in $L=\mathbb{R}^{d}$ with $f$ being indicator that the distance is at most some threshold. If $f,L$ is fixed (i.e., independent of the number of rows/columns) then these graphs are closed under row/column deletion and thus have bounded VC-dimension by Theorem˜1.1.
@@
Shortest-path structures. Let $G$ be an undirected graph with non-negative edge weights and let $S$ be a subset of its shortest paths such that, for every pair $(u,v)$ of distinct vertices, $S$ contains exactly one shortest path between $u$ and $v$. Marie de Lima et al. [2023] defines a range space associated with $S$ and proves that its VC dimension is 2.
@@
Adjacency matrices of planar graphs. Let $G$ be a planar graph. Then, there is a planar drawing of $G$, and deleting vertices and edges retains this property. Therefore, these graphs are closed under row/column deletion and thus have bounded VC-dimension by Theorem˜1.1.
@@
Adjacency matrices of semi-algebraic graphs of bounded description complexity. By the Milnor-Thom theorem in real algebraic geometry [Matoušek, 2002] and Assouad’s theorem [Assouad, 1983], semi-algebraic graphs of bounded description complexity have constant VC-dimensions. In turn, this extends to adjacency matrices of string graphs where every two curves intersect.
@@
Practical matrices. Coudert et al. [2024] computes the VC-dimension of families of graphs that arise in practice, from protein interaction networks to autonomous Internet systems. Even for such graphs with millions of nodes, the VC-dimension of the graphs are typically between $3$ and $8$.
Appendix E Numerical Simulations
The algorithm from Theorem˜1.3 has previously been described in Björklund and Lingas [2001] and Alves et al. [2024]. The latter already experimentally verified the efficiency of the algorithm on real-world data sets and observed that it beats the naive quadratic time matrix vector multiplication. Our work proves theoretic guarantees for this speed-up parameterized by the (corrupted) VC-dimension. While the main contribution of this work is to have theoretic guarantees and explanation for why the algorithm is efficient in practice, we here complement the result with a few experiments [^1] The code for the experiments is available at https://github.com/emiletimothy/structural_complexity_of_matrix_vector_multiplication. As we give theoretic bounds parameterized by $d$, our experiments focus on the $d$-dependence too.
Matrix Generation. Since we want to study the complexity dependence on $d$, we fix the matrix dimension to $n=2^{12}=4096$ and run experiments for d=2,,12d=2,...,12. The maximum $d$ is 12 because the VC-dimension $\leq\log(n)=12$ so we do not need to consider $d>\log(n)$. The matrix size $n$ was picked due to execution time constraints. The $n\times n$ Binary matrices are generated by sampling i.i.d uniform $a^{(i)}\in[0,1]^{d},b^{(i)}\in[0,d/2]$ for i=1,,ni=1,...,n and $x^{(j)}\in[0,1]^{d}$ for j=1,,nj=1,...,n, and letting $\mathbf{M}_{i,j}=\mathbbm{1}_{\{a^{(i)\top}x^{(j)}\geq b^{(i)}\}}$. The range for the $b^{(i)}$ was chosen so that neither $\mathbf{M}$ nor $1-\mathbf{M}$ are sparse, which otherwise would allow for naive matrix multiplication to already beat $O(n^{2})$ time. The set system of linear classifiers has VC-dimension $d$, so the corrupted VC-dimension of $\mathbf{M}$ is upper bounded by $d$. This implies that the respective MST has weight at most $\widetilde{O}(n^{2-1/d})$. Figure˜1 shows the observed weight of the MST relative to the ambient dimension $d$. The MST was computed exactly since easier to implement and the preprocessing time complexity is not focus of this work.
Complexity. For the matrix vector products we generate i.i.d. uniform vectors $v\in[0,1]^{n}$. For each $d$, we compute $1000$ matrix vector products and measure the run time. Figure˜2 shows the mean and standard deviation of the observed time. We also compare the time to the naive numpy matrix vector multiplication. Figure˜2 shows that, since matrix size $n$ is fixed, the time complexity of numpy is constant for different $d$, whereas the MST-based algorithm is substantially faster for small $d$. As $d$ increases, the gap gets smaller as one would expect, given the $\widetilde{O}(n^{2-1/d})$ theoretic upper bound. The experiments were conducted on an Apple M4 Pro macOS model with 12 physical cores / threads, and 48 GB unified memory RAM, and an Apple Accelerate Framework BLAS backend. The total amount of time to run the experiments was $\approx 30$ minutes. No GPU acceleration was used; all matrix operations ran on CPU.
[…]
Figure 1: Weight of the $\Delta$-labeled spanning tree for $2^{12}\times 2^{12}$ matrices generated from halfspaces in $d$ dimensions.
224466881010121200.20.20.40.40.60.60.80.8Ambient Dimension of Set SystemRuntime (ms)
Figure 2: Average run time of the MST-based matrix vector algorithm (bold) and naive numpy matrix vector multiplication (blank), with 1-sigma error bars. Horizontal axis is the dimension of the points and halfspaces used to construct the matrix.